OpenStudy (anonymous):

There are 100 students in South High School. South High School offers only Chinese and Spanish. There are 8 more students in Chinese than in Spanish, and every student takes at least one language. If 37 students take only Spanish, then how many take both languages?

OpenStudy (anonymous):

The equation is 8+S+S=100 2S+8=100 2S=100-8 S=92/2 S=46 46-37=9